# Multilayer Artificial Benchmark for Community Detection (mABCD)

### v1

@@ 3.1 Correlations Between Layers
Although there are no edges between nodes in different layers, in most real-world multilayered networks, layers are clearly not independently generated. Each actor is associated with $\ell$ nodes, one in each layer, and there are some highly non-trivial correlations between edges across layers. For example, active users on one social media platform are often also active on another one [27]. This creates correlations between degree distributions across layers. Communities that are naturally formed in various layers often depend on the properties of the associated actors. For example, users interested in soccer might group together on Instagram and on Facebook. As a result, partitions of nodes into communities (associated with different layers) are often correlated. Finally, interactions between actors in one layer might increase their chances of interacting in another layer, yielding correlations at the level of edges.
Below, we briefly summarize how we measure these three types of correlations mentioned above. The first two measures are standard, and their detailed description can be found, for example, in [12, 35].
Correlations Between Nodes Degrees in Various Layers
We will use Kendall rank correlation coefficient $\tau$ [39] to measure correlations between sequences of node degrees in two different layers. It is a nonparametric measure of the ordinal association between two measured quantities: the similarity of the orderings of the data when ranked by each of the quantities (in our application, the degree sequences). The Kendall correlation between two variables ranges from $-1$ to $1$. It is large when observations have a similar rank between the two variables and is small when observations have a dissimilar rank between the two variables.
@@
The adjusted mutual information (AMI), a variation of mutual information (MI), is a common way to compare partitions of the same set [35, 55]. Usually, one may want to compare the partitions returned by some clustering algorithms. In our present context, we may want to compare partitions into ground-truth communities from two different layers. The AMI takes a value of $1$ when the two partitions are identical and 00 when the MI between two partitions equals the value expected due to chance alone. Actors, that are inactive in at least one of the two layers we compare against each other, are ignored so that a comparison of partitions is made on the same set of actors.
@@
$E_{i}^{j}=\{a_{1}a_{2}:(a_{1},i)(a_{2},i)\in E\wedge a_{1},a_{2}\in[n]\wedge a_{1}\text{ and }a_{2}\text{ are active in layers $i$ and $j$}\},$ (1)
@@
Note that the definition of $\mathbf{R}$ implies that $r_{i,i}=1$ for any $i\in[\ell]$ and $r_{i,j}=r_{j,i}$ for $1\leq i<j\leq\ell$. The maximum value of $1$ is attained when edges in one of the layers form a subset of edges in the other layer. The minimum value of 00 is attained when the two sets of edges in the corresponding layers are completely disjoint. As a result, $r_{ij}$ aims to capture correlations between individual edges, but it is not normalized as, for example, the Kendall rank correlation coefficient $\tau$. The coefficient $\tau$ ranges from $-1$ to $1$, corresponding to the two extremes, and 00 corresponds to a neutral case. Graphs associated with the layers are sparse, but one layer might have substantially more edges than the other. Hence, $r_{ij}$ is convenient, but it does not have a natural interpretation as $\tau$. Finally, let us mention that one can easily update the value of $r_{ij}$ when some small operations are applied to either layer $i$ or $j$. It will become handy when such operations must be performed on our synthetic model to converge to the desired correlation matrix $\mathbf{R}$.
@@ 3.2 Examples of Multilayer Networks
arxiv 13 14,065 26,796 59,026 Coauthorship network obtained from articles published on the “arXiv” repository [20]. Each layer represents a different arXiv category, e.g. Physics and Society or Social and Information Networks.
aucs 5 61 224 620 A graph of interactions (friends on Facebook, leisure, work, co-authorship and lunch) between employees of Aarhus University, Department of Computer Science [50].
cannes 3 438,537 659,951 974,743 A network of interactions (retweets, mentions and replies) between Twitter (now X) users during the Cannes Film Festival in 2013 [45].
ckmp 3 241 674 1,370 A network depicting diffusion of innovations among physicians [17]. Each layer was built based on the physician’s answers to questions like “Tell me the first names of your three friends whom you see most often socially”.
eutr-A 37 417 2,034 3,588 The European air transportation network [15]. Each layer represents the connections of different airline operator in Europe.
l2-course 2 41 82 297 A network of interactions (collaboration, friends) between U.S. students learning Arabic language during an intensive course in Jordan (1st month snapshot) [48].
lazega 3 71 212 1,659 A network of interactions (co-work, friendship and advice) between staff of a law corporation [51].
timik 3 61,702 102,247 881,676 A graph of interactions (text messages, online transactions, home visits) between users of the virtual world platform for teenagers [32].
Next, in Figure 2 we present correlation matrices between layers for degrees of nodes, partitions, and edges, according to the methodology desctibed in Section 3.1. Due to a lack of ground truth partitions, communities were identified by the outcome of the Louvain algorithm [9]. After conducting a thorough analysis of the results, we observed a lack of consistent patterns in the examined real-world multilayer networks. Specifically, there was no universal relationship between degrees, edges, and community partitions across the layers. In some cases, these features were correlated, while in others, they appeared to be independent. Furthermore, within a single network, we found that while two layers might exhibit a correlation, this relationship did not necessarily extend to other layers. For instance, a node might have similar neighbours across two layers but exhibit a completely different connectivity pattern in a third layer. These observations highlight the inherent heterogeneity and complexity of real-world multilayer networks, emphasizing the need for a flexible and customizable synthetic network generation framework.
@@ 3.3 Assumed Functionalities
In light of these findings, we assumed that the framework has to accommodate inter-layer diversities. Moreover, since the question of how precisely these relations should be modelled constitutes an issue that must be addressed individually, depending on the specific problem one is working on, we aimed to provide a tool that balances flexibility and accessibility. Thus, building the framework upon ABCD seemed a plausible approach, as the design assumptions underlying it allow for extension in this direction.
@@ 4 The mABCD Model
In this section, we introduce a variant of the ABCD model that produces a synthetic collection of graphs that form a multilayer structure, mABCD.
4.1 Power-law Distribution
Power-law distributions will be used to generate both the degree sequence and community sizes so let us formally define it. For given parameters $\gamma\in(0,\infty)$, $\delta,\Delta\in{\mathbb{N}}$ with $\delta\leq\Delta$, we define a truncated power-law distribution $\mathcal{P}\left(\gamma,\delta,\Delta\right)$ as follows. For $X\sim\mathcal{P}\left(\gamma,\delta,\Delta\right)$ and for $k\in{\mathbb{N}}$ with $\delta\leq k\leq\Delta$,
[…]
4.2 The Configuration Model
The well-known configuration model is an important ingredient of the generation process, so let us formally define it here. Suppose then that our goal is to create a graph on $n$ nodes with a given degree sequence $\textbf{d}=(d_{i},i\in[n])$, where d is a sequence of non-negative integers such that $m=\sum_{i\in[n]}d_{i}$ is even. We define a random multi-graph $\mathrm{CM}(\textbf{d})$ with a given degree sequence known as the configuration model (sometimes called the pairing model), which was first introduced by Bollobás [10]. (See [7, 56, 57] for related models and results.)
@@ 4.3 Parameters of the mABCD Model
The mABCD model is governed by the following parameters. The first family of parameters is responsible for a few global properties of the model and is enlisted in Table 2.
@@
[…]
$\mathbf{R}$ $[0,1]^{\ell\times\ell}$ Correlation between edges
[…]
@@
The second family of parameters (Table 3) is responsible for various properties that are specific for each of the $\ell$ layers; subscripts $i\in[\ell]$ indicate that the corresponding parameters shape the $i^{th}$ layer. In particular, the set of parameters $\xi_{i}$, $i\in[\ell]$, will control the level of noise, that is, the fraction of edges in layer $i$ that are between nodes from two different communities.
Table 3: Layer-specific parameters of mABCD.
@@
$\tau_{i}$ $[-1,1]$ Correlation coefficient between degrees and labels
@@
[…]
$\Delta_{i}$ ${\mathbb{N}}\;(1\leq\delta_{i}\leq\Delta_{i}<n)$ Max degree at most $\Delta_{i}$
@@
The suggested range of values for parameters $\gamma$ and $\beta$ (namely, the intervals $(2,3)$ and $(1,2)$, respectively) are chosen according to experimental values commonly observed in complex networks [3, 46].
@@ 4.4 The mABCD Construction
The degree sequences for all of the $\ell$ layers are generated independently so we may concentrate on a given layer $i\in[\ell]$. We ensure that the degree sequence satisfies (a) a power-law with parameter $\gamma_{i}$, (b) a minimum value of at least $\delta_{i}$, and (c) a maximum value of at most $\Delta_{i}$.
@@
Similarly to the process of generating the degree sequences, the sequence of community sizes are generated independently, ensuring that the distribution for a given layer $i\in[\ell]$, satisfy (a) a power-law with parameter $\beta_{i}$, (b) a minimum value of $s_{i}$, and (c) a maximum value of $S_{i}$. In addition, we also require that the sum of community sizes is exactly $n$. Specifically, inactive nodes (if there are any) form their own community, namely, $C^{i}_{0}$. Other communities are generated with sizes determined independently by the distribution $\mathcal{P}\left(\beta_{i},s_{i},S_{i}\right)$. We generate communities until their collective size is at least $n$. If the sum of community sizes at this moment is $n+x$ with $x>0$, then we perform one of two actions: if the last added community has a size at least $x+s_{i}$, then we reduce its size by $x$. Otherwise (that is, if its size is $c<x+s_{i}$), then we delete this community, select $c-x$ old communities and increase their sizes by 1.
[…]
@@
The above strategy creates a partition of nodes that is highly correlated with the geometric locations of nodes in the reference layer; nodes that are close to each other in the reference layer are often in the same community — see Figure 3 for an example. To reduce the correlation strength (modelled by the parameter $r_{i}\in[0,1]$), we perform the following procedure. Each active node independently leaves its own community with probability $1-r_{i}$, freeing a spot in this community. All the nodes that left are then put back randomly to any available spot (which typically is in a community that this node was not originally in). Note that in the extreme case, when $r_{i}=0$, the resulting partition does not depend on the reference layer at all, so there is no correlation. We write $\textbf{C}^{i}_{n}=(C^{i}_{j},j\in\{0\}\cup[L^{i}])$ for the generated collection of communities in $G^{i}_{n}$. Again, let us stress the fact that $\textbf{C}^{i}_{n}$ is a random partition of $[n]$ of random size $L^{i}+1$.
@@
Figure 3: Two partitions generated based on the same reference layer with $n=1{,}000$ nodes: (left) $q_{1}=1$ (all nodes active), $S_{1}=32$, $s_{1}=16$, $\beta_{1}=1.5$, (right): $q_{2}=0.5$ (50% nodes active), $S_{2}=50$, $s_{2}=25$, $\beta_{2}=1.5$.
@@
Now, it is time to form edges in mABCD. It will be done in the next three phases, Phases 4–6. Phases 4 and 5 will independently generate $\ell$ graphs $G^{i}_{n}$, $i\in[\ell]$, for each of the $\ell$ layers whereas Phase 6 will make sure that edges across various layers are correlated, if needed. We may then concentrate on a given layer $i\in[\ell]$.
@@
[…]
@@
The algorithm is implemented in Julia programming language, a high-level, general-purpose dynamic programming language, designed to be fast and productive. Source code and installation instructions are available on mABCD GitHub repository [^5] https://github.com/KrainskiL/MLNABCDGraphGenerator.jl. The Python ports are available as well [^6] https://github.com/anty-filidor/spreading-vs-mln-structure.

### v2

@@ 3.1 Correlations between layers
Although there are no edges between nodes in different layers, in most real-world multilayered networks, layers are clearly not independently generated. Each actor is associated with $\ell$ nodes, one in each layer, and there are some highly non-trivial correlations between edges across layers. For example, active users on one social media platform are often also active on another one [gottfried2024americans]. This creates correlations between degree distributions across layers. Communities that are naturally formed in various layers often depend on the properties of the associated actors. For example, users interested in soccer might group together on Instagram and on Facebook. As a result, partitions of nodes into communities (associated with different layers) are often correlated. Finally, interactions between actors in one layer might increase their chances of interacting in another layer, yielding correlations at the level of edges.
Below, we summarize how we measure these three types of correlations. The first two measures are standard, and their description can be found in [brodka2018quantifying, kaminski2021mining].
Correlations between node degrees in various layers
We will use Kendall rank correlation coefficient $\tau$ [kendall1938new] to measure correlations between sequences of node degrees in two different layers. It is a nonparametric measure of the ordinal association between two measured quantities: the similarity of the orderings of the data when ranked by each of the quantities (in our application, the degree sequences). The Kendall correlation between two variables ranges from 1-1 to 11. It is large when observations have a similar rank between the two variables and is small when observations have a dissimilar rank.
@@
The adjusted mutual information (AMI), a variation of mutual information (MI), is a common way to compare partitions of the same set [kaminski2021mining, vinh2009information]. Usually, one may want to compare the partitions returned by some clustering algorithms. In our present context, we may want to compare partitions into ground-truth communities from two different layers. The AMI takes a value of $1$ when the two partitions are identical and 0 when the MI between two partitions equals the value expected due to chance alone. Actors, that are inactive in at least one of the two layers we compare against each other, are ignored so that a comparison of partitions is made on the same set of actors.
@@
[…]
@@
Note that the definition of $\mathbf{R}$ implies that $r_{i,i}=1$ for any $i\in[\ell]$ and $r_{i,j}=r_{j,i}$ for $1\leq i<j\leq\ell$. The maximum value of $1$ is attained when edges in one of the layers form a subset of edges in the other layer. The minimum value of 0 is attained when the two sets of edges in the corresponding layers are completely disjoint. As a result, $r_{ij}$ aims to capture correlations between individual edges, but it is not normalized as, for example, the Kendall rank correlation coefficient $\tau$. The coefficient $\tau$ ranges from 1-1 to $1$, corresponding to the two extremes, and 0 corresponds to a neutral case. Graphs associated with the layers are sparse, but one layer might have substantially more edges than the other. Hence, $r_{ij}$ is convenient, but it does not have a natural interpretation as $\tau$. Finally, let us mention that one can easily update the value of $r_{ij}$ when some small operations are applied to either layer $i$ or $j$. It will become handy when such operations must be performed on our synthetic model to converge to the desired correlation matrix $\mathbf{R}$.
@@ 3.2 Examples of multilayer networks
arxiv 13 14,065 26,796 59,026 Coauthorship network obtained from articles published on the “arXiv” repository [dedomenico2015arxiv]. Each layer represents a different arXiv category, e.g. Physics and Society or Social and Information Networks.
aucs 5 61 224 620 A graph of interactions (friends on Facebook, leisure, work, co-authorship and lunch) between employees of Aarhus University, Department of Computer Science [rossi2015aucs].
cannes 3 438,537 659,951 974,743 A network of interactions (retweets, mentions and replies) between Twitter (now X) users during the Cannes Film Festival in 2013 [omodei2015characterizing].
ckmp 3 241 674 1,370 A network depicting diffusion of innovations among physicians [coleman1957ckmp]. Each layer was built based on the physician’s answers to questions like “Tell me the first names of your three friends whom you see most often socially".
eutr-A 37 417 2,034 3,588 The European air transportation network [cardillo2013eutransportation]. Each layer represents the connections of different airline operator in Europe.
l2-course 2 41 82 297 A network of interactions (collaboration, friends) between U.S. students learning Arabic language during an intensive course in Jordan (1st month snapshot) [paradowski2024peer].
lazega 3 71 212 1,659 A network of interactions (co-work, friendship and advice) between staff of a law corporation [snijders2006lazega].
timik 3 61,702 102,247 881,676 A graph of interactions (text messages, online transactions, home visits) between users of the virtual world platform for teenagers [jankowski2017timik].
Next, in Figure 2, we present correlation matrices between layers for degrees of nodes, partitions, and edges, according to the methodology described in Section 3.1. Due to a lack of ground truth partitions, communities were identified by the outcome of the Louvain algorithm [blondel2008fast]. After conducting a thorough analysis of the results, we observed a lack of consistent patterns in the examined real-world multilayer networks. Specifically, there was no universal relationship between degrees, edges, and community partitions across the layers. In some cases, these features were correlated, while in others, they appeared to be independent. Furthermore, within a single network, we found that while two layers might exhibit a correlation, this relationship did not necessarily extend to other layers. For instance, a node might have similar neighbours across two layers but exhibit a completely different connectivity pattern in a third layer. These observations highlight the inherent heterogeneity and complexity of real-world multilayer networks, emphasizing the need for a flexible and customizable synthetic network generation framework.
@@ 3.3 Assumed functionalities
In light of these findings, we assumed that the framework has to accommodate inter-layer diversities. Moreover, since the question of how precisely these relations should be modelled constitutes an issue that must be addressed individually, depending on the specific problem one is working on, we aimed to provide a tool that balances flexibility and accessibility. Thus, building the framework upon ABCD seemed a plausible approach, as it allows for extension in this direction.
@@ 4 The mABCD Model
In this section, we introduce a variant of the ABCD model that produces a synthetic collection of graphs that form a multilayer structure, mABCD. The subsequent text guides through the details of the network construction process. Its parameters are summarized in Table 2 and Table 3, while the process itself is is conceptually illustrated in Figure 3.
4.1 Preliminaries
Since the networks produced by the proposed framework follow a power-law distribution in both node degrees and community sizes, and since their generation proceeds sequentially, we first introduce the preliminary concepts that are fundamental to the construction process.
Power-law distribution
Power-law distributions will be used to generate both the degree sequence and community sizes. For given parameters $\gamma\in(0,\infty)$, $\delta,\Delta\in{\mathbb{N}}$ with $\delta\leq\Delta$, we define a truncated power-law distribution $\mathcal{P}\left(\gamma,\delta,\Delta\right)$ as follows. For $X\sim\mathcal{P}\left(\gamma,\delta,\Delta\right)$ and for $k\in{\mathbb{N}}$ with $\delta\leq k\leq\Delta$,
$\mathbb{P}\left(X=k\right)=\frac{\int_{k}^{k+1}x^{-\gamma}\,dx}{\int_{\delta}^{\Delta+1}x^{-\gamma}\,dx}\,.$
The configuration model
Suppose then that our goal is to create a graph on $n$ nodes with a given degree sequence $\textbf{d}=(d_{i},i\in[n])$, where d is a sequence of non-negative integers such that $m=\sum_{i\in[n]}d_{i}$ is even. We define a random multi-graph $\mathrm{CM}(\textbf{d})$ with a given degree sequence known as the configuration model (sometimes called the pairing model), which was first introduced by [bollobas1980probabilistic]. See: [bender1978asymptotic, wormald1984generating, wormald1999models] for related models and results.
@@ 4.2 Parameters of the mABCD model
The mABCD model is governed by the following parameters. The first family of parameters is responsible for a few global properties of the model and is listed in Table 2.
@@
$n$ $\mathbb{N}$} Number of actors (recommended $n\geq 1{,}000$)
$\ell$ $\mathbb{N}$} Number of layers
$\mathbf{R}$ $0,1]^{\ell\times\ell}$ Inter-layer edge correlation matrix
$d$ $\mathbb{N}$} Dimension of reference layer
@@
The second family of parameters (Table 3) is responsible for various properties that are specific to each of the $\ell$ layers; subscripts $i\in[\ell]$ indicate that the corresponding parameters shape the $i^{th}$ layer. In particular, the set of parameters $\xi_{i}$, $i\in[\ell]$, will control the level of noise, that is, the fraction of edges in layer $i$ that are between nodes from two different communities. The suggested range of values for parameters $\gamma$ and $\beta$ (namely, the intervals $(2,3)$ and $(1,2)$, respectively) are chosen according to experimental values commonly observed in complex networks [barabasi2016network, orman2009comparison].
Table 3: Layer-specific parameters of mABCD (the index denotes the $i$-th layer).
@@
$\tau_{i}$ [1,1][-1,1] Correlation coefficient between degrees and labels
@@
$\delta_{i}$ $\mathbb{N}$} Min. degree at least $\delta_{i}$ (recommended $\delta_{i}\geq 10$)
$\Delta_{i}$ ${\mathbb{N}}\;(1\leq\delta_{i}\leq\Delta_{i}<n)$ Max. degree at most $\Delta_{i}$
@@ 4.3 The mABCD construction
Conceptually, the phases of network construction are illustrated in Figure 3. The first five steps are carried out independently for each of the $\ell$ layers, while the sixth binds the generated graphs into a single multilayer network. The process is governed by the parameters listed in Table 2 and Table 3. In the first step, the so-called active nodes (i.e., those that will not be isolated) are determined. Each active node is then assigned its degree value. During the third step, communities are created and populated with nodes according to the latent, biscuit-like layer. Phase four focuses on connecting nodes within the same community as well as establishing inter-community (i.e. background) links. These graphs are subsequently simplified to remove self-loops and multiedges while preserving the intended structural properties. The final phase performs a series of edge rewirings to achieve the targeted edge correlations between each pair of layers. The illustration is complemented by pseudocode published on GitHub repository.
Figure 3: A schematic illustration of the mABCD model’s operating principle, exemplified by the generation of a three-layer network. The process consists of six steps. The first five are performed independently for each layer, yielding scale-free monoplex graphs with a known community structure drawn from a shared latent space. In the final step, the layers are glued through a sequence of rewirings to achieve the desired cross-layer edge correlations.
@@
The degree sequences for all of the $\ell$ layers are generated independently so we may concentrate on a given layer $i\in[\ell]$. We ensure that the degree sequence satisfies (a) a power-law with parameter $\gamma_{i}$, (b) a minimum value of at least $\delta_{i}$, and (c) a maximum value of at most $\Delta_{i}$ (see Section 4.1).
@@
Similarly to the process of generating the degree sequences, the sequence of community sizes are generated independently, ensuring that the distribution for a given layer $i\in[\ell]$, satisfy (a) a power-law with parameter $\beta_{i}$, (b) a minimum value of $s_{i}$, and (c) a maximum value of $S_{i}$ (see Section 4.1). In addition, we also require that the sum of community sizes is exactly $n$. Specifically, inactive nodes (if there are any) form their own community, namely, $C^{i}_{0}$. Other communities are generated with sizes determined independently by the distribution $\mathcal{P}\left(\beta_{i},s_{i},S_{i}\right)$. We generate communities until their collective size is at least $n$. If the sum of community sizes at this moment is $n+x$ with $x>0$, then we perform one of two actions. If the last added community has a size at least $x+s_{i}$, then we reduce its size by $x$. Otherwise (that is, if its size is $c<x+s_{i}$), we delete this community, select cxc-x old communities and increase their sizes by 1.
Now, given that the sequences of community sizes are already determined (for all layers), it is time to assign nodes to communities. To allow communities to be correlated with each other, we first create a latent reference layer that will guide the process of assigning nodes to specific communities across all layers. One may think of this auxiliary layer as properties of actors (such as people’s age, education, geographic location, beliefs, etc.) shaping different layers (for example, various social media platforms). This single reference layer will be used for all $\ell$ layers. In this reference layer, each actor $a\in[n]$ gets assigned a random vector in $\mathbb{R}^{d}$ (by default, $d=2$) that is taken independently and uniformly at random from the ball of radius one centred at $\textbf{0}=(0,0,\ldots,0)$.
@@
The above strategy creates a partition of nodes that is highly correlated with the geometric locations of nodes in the reference layer; nodes that are close to each other in the reference layer are often in the same community. An example of this process is illustrated in Figure 4, which shows how nodes are assigned to communities in two network layers, both guided by the same reference layer. This mechanism can be intuitively compared to eating a biscuit, bite by bite, from the edge towards the centre. Nodes that are close to each other are more likely to be “eaten at the same time” and thus placed in the same group. Moreover, the severed pieces may vary in size, so do communities. Finally, the same biscuit can be enjoyed by two different gourmets, which in this analogy corresponds to reusing the same reference layer for each of the relationships modelled within the generated network. To reduce the correlation strength (modelled by the parameter $r_{i}\in[0,1]$), we perform the following procedure. Each active node independently leaves its own community with probability $1-r_{i}$, freeing a spot in this community. All the nodes that left are then put back randomly to any available spot (which typically is in a community that this node was not originally in). Note that in the extreme case, when $r_{i}=0$, the resulting partition does not depend on the reference layer at all, so there is no correlation. We write $\textbf{C}^{i}_{n}=(C^{i}_{j},j\in\{0\}\cup[L^{i}])$ for the generated collection of communities in $G^{i}_{n}$. Again, let us stress the fact that $\textbf{C}^{i}_{n}$ is a random partition of $[n]$ of random size $L^{i}+1$. At this point, it is worth mentioning that, due to the shared latent layer and a parameter $r_{i}$ specific to each relationship, it is possible to model the alignment of communities across the network. Following [magnani2021community], where the notion of a pillar is introduced to describe a community that persists across at least two layers, layers with $r_{i}\approx 1$ tend to exhibit this form of cluster alignment. Conversely, if the values of $r_{i}$ are set to more distinctive numbers, the communities are expected to be largely disjoint across the layers.
Figure 4: Two partitions generated based on the same reference layer with n=1,000n=1{,}000 nodes: (left) $q_{1}=1$ (all nodes active), $S_{1}=32$, $s_{1}=16$, $\beta_{1}=1.5$, (right): $q_{2}=0.5$ (50% nodes active), $S_{2}=50$, $s_{2}=25$, $\beta_{2}=1.5$ with inactive nodes shown in grey.
@@
Now, it is time to form edges in mABCD. It will be done in the next three phases. Phases 4 and 5 will independently generate $\ell$ graphs $G^{i}_{n}$, $i\in[\ell]$, for each of the $\ell$ layers whereas Phase 6 will make sure that edges across various layers are correlated, if needed. We may then concentrate on a given layer $i\in[\ell]$.
@@
(Note that ${\mathbb{E}}\left[\left\lfloor a+b\right\rceil\right]=a(1-b)+(a+1)b=a+b$.) Now define $Y_{a}=\left\lfloor(1-\xi)\deg^{i}(a)\right\rceil$ and $Z_{a}=\deg^{i}(a)-Y_{a}$ (note that $Y_{a}$ and $Z_{a}$ are random variables with ${\mathbb{E}}\left[Y_{a}\right]=(1-\xi)\deg^{i}(a)$ and ${\mathbb{E}}\left[Z_{a}\right]=\xi\deg^{i}(a)$ and since we generate each layer separately they are different for each layer) and, for all $a\in[n]$, split the $\deg^{i}(a)$ half-edges of $(a,i)$ into $Y_{a}$ community half-edges and $Z_{a}$ background half-edges. Next, for all $j\in[L^{i}]$, construct the community graph $G^{i}_{n,j}$ as per the configuration model (see Section 4.1) on node set $C^{i}_{j}$ and degree sequence $(Y_{a},a\in C^{i}_{j})$. Note that $C_{0}$ consists of inactive nodes which, by design, have degree zero. Hence, there is no need to do anything with them. Finally, construct the background graph $G^{i}_{n,0}$ as per the configuration model on node set $[n]$ and degree sequence $(Z_{a},a\in[n])$. In the event that the sum of degrees in a community is odd, we pick a maximum degree node $(a,i)$ in said community and replace $Y_{a}$ with $Y_{a}+1$ and $Z_{a}$ with $Z_{a}-1$. Note that $G^{i}_{n,j}$ is a graph, and $C^{i}_{j}$ is the set of nodes in this graph; we refer to $C^{i}_{j}$ as a community and $G^{i}_{n,j}$ as a community graph. Note also that $G^{i}_{n}=\bigcup_{0\leq j\leq L^{i}}G^{i}_{n,j}$.
@@
The algorithm is implemented in Julia programming language, a high-level, general-purpose dynamic programming language, designed to be fast and effective. Source code and documentation are available on mABCD GitHub repository [^6] https://github.com/KrainskiL/MLNABCDGraphGenerator.jl. The Python ports are available as well [^7] https://github.com/anty-filidor/spreading-vs-mln-structure.